I have a word problem that i can't figure out would you please help me.
Question- write a exponential function to model the situation. Then estimate the value of the function after 5 years(to the nearest whole number)-- A population of 270 animals that decreses at an anual rate of 22%

I am not totaly sure what formula to use? I think it might be one of the following. A=p(1+r/n)or A=p(1-r/n) Im not sure. Could you please help me to decide what formula to use. Thank you very much!!!

Answer by stanbon(57219)

Question- write a exponential function to model the situation. Then estimate the value of the function after 5 years(to the nearest whole number)-- A population of 270 animals that decreses at an anual rate of 22%------------------------------------------
f(t) = 270*0.78^(t)
f(5) = 270*0.78^5
f(5)= 270*0.28872
f(5)= 77.95
since you cannot have 0.95 or an animal f(5) = 77 animals after 5 years
Cheers,
Stan H.